OpenStudy (firejay5):

A 1200kg car is traveling at 25m/s. How much kinetic energy does the car have? What amount of work would be required to bring the car to a stop?

OpenStudy (firejay5):

KE = 1/2mv^2 = (0.5)(1200)(25)^2 = 375,000 J

OpenStudy (anonymous):

Assuming that the car is travelling in a horizontal plane on a frictionless surface, the work-kinetic energy theorem shows that it would take 375,000 J of work to bring the car to a stop.

OpenStudy (firejay5):

@ospreytriple How would I show work for that

OpenStudy (firejay5):

It said W = KEf - KEi

OpenStudy (firejay5):

It said W = KEf - KEi

OpenStudy (anonymous):

Try this; \[W=\Delta E_k\]\[=E_{k,f} - E_{k,i}\]\[=0-375,000 \text{ J}\]\[=-375,000 \text{ J}\] and since energy is being removed from the object, work is negative.